Given P(A) = .70, P(B) = .30, and P(A ∩ B) = .00, find (a) P(A ∪ B) and (b) P(A | B).
Answer by stanbon(48546)

Given
P(A) = .70,
P(B) = .30, and
P(A ∩ B) = .00,
find
(a) P(A ∪ B)
P(A U B) = P(A)+P(B)-P(AnB)
= 0.7+0.3 + 0 = 1.
---------------------------
(b) P(A | B) = P(AnB)/P(B) = 0/0.3 = 0
=============================================
Cheers,
Stan H.